### v1

# Dual Numbers for Arbitrary Order Automatic Differentiation

Abstract
Dual numbers are a well-known tool for computing derivatives of functions. While the theoretical framework for calculating derivatives of arbitrary order is well established, practical implementations remain less developed. One notable implementation is available in the Julia programming language where dual numbers are designed to be nested, enabling the computation of derivatives to arbitrary order. However, this approach has a significant drawback as it struggles with scalability for high-order derivatives. The nested structure quickly consumes memory, making it challenging to compute derivatives of higher orders. In this study, we introduce DNAOAD, a Fortran-based implementation of automatic differentiation capable of handling derivatives of arbitrary order using dual numbers. This implementation employs a direct approach to represent dual numbers without relying on recursive or nested structures. As a result, DNAOAD facilitates the efficient computation of derivatives of very high orders while addressing the memory limitations of existing methods.
Facultad de Ingeniería, Universidad Autónoma de Yucatán, A.P. 150, Cordemex, Mérida, Yucatán, México.
Keywords: Dual numbers, Automatic Differentiation, Fortran.
1 Introduction
Differentiation is a fundamental process that is widely applied in numerous fields of science and engineering. It can be carried out symbolically using specialized software, such as Maxima, Maple, or Mathematica, or numerically through finite-difference methods. Another approach is Automatic Differentiation (AD) [1, 2, 3], which enables the numerical computation of derivatives with the same level of accuracy and precision as symbolic methods, but without requiring symbolic computations.
Automatic Differentiation (AD) is a systematic algorithmic technique for efficiently applying the chain rule to compute derivatives. Although there is extensive research on this topic, most studies focus on the computation of first- and second-order derivatives for real-valued functions [4, 5, 6, 7, 8, 9, 10]. ADOL-C [11, 12, 13] is an interesting implementation of this technique, which takes advantage of operator overloading to enable differentiation. Although ADOL-C supports higher-order derivatives for some functions, a considerable number, including asin, acos, atan, and exponentiation, lack full support for differentiation across all derivative orders.
Notable AD implementations that support the computation of derivatives of arbitrary order include the PyTorch library in Python [14] and the ForwardDiff package in Julia [15]. Even when these libraries do not natively support automatic differentiation for complex-valued functions, the complex case can be managed by separately computing the derivatives of the real and imaginary components of the functions involved.
PyTorch is a versatile and widely used Python library for AD and deep learning applications. One of its key strengths is its ability to compute higher-order derivatives using a dynamic computation graph. This graph tracks operations at runtime and facilitates automatic computation of gradients. However, when dealing with higher-order differentiation—especially when calculating derivatives of complex operations or multiple gradients—PyTorch retains the entire computational graph in memory. While this is necessary to compute successive gradients, it significantly increases memory consumption. For instance, calculating higher-order derivatives can cause the computational graph to rapidly expand due to the accumulation of intermediate values. Consequently, PyTorch’s memory usage may quickly exceed the system’s available memory, often resulting in a std::bad_alloc error, which occurs when memory allocation requests surpass the system’s capacity. In our numerical experiment, performed on a 12th generation Intel(R) Core(TM) i7-12650H CPU @ 4.70 GHz (base 400 MHz) with 15 GB of RAM, running Debian GNU/Linux 12, this error occurred when calculating the 12th-order derivative of the function $f(x)=\sin(x)\,\exp(-x^{2})$, composed with itself five times.
In contrast, ForwardDiff, implemented in Julia, computes derivatives using dual numbers and elegantly handles higher-order derivatives through their recursive nesting. For simplicity, we refer to a structure resulting from nesting these numbers $n$ times as a dual number of order $n$ or a multidual number of order $n$ [16, 17, 18, 19]. Despite the elegance of this approach, challenges emerge when computing higher-order derivatives with ForwardDiff. With each successive derivative, ForwardDiff generates a new Dual type that encapsulates all previously computed derivative information. This recursive nesting increases complexity, as each successive derivative introduces additional layers of dual numbers. For higher-order derivatives, this nesting becomes particularly problematic, leading to excessive memory consumption and ultimately causing stack overflows. In our numerical experiments, a StackOverflowError occurred when attempting to compute the 15th-order derivative of the previously mentioned function.
In this study, we present a Fortran implementation of dual numbers to arbitrary order without relying on a recursive or nested structure. This approach theoretically enables the computation of derivatives of any order, effectively overcoming the limitations of the nested dual number method. While practical constraints, such as computational resources, will ultimately limit the achievable order of derivatives, our method significantly extends the range of computable derivatives beyond what is feasible with the nested approach. It is important to note, however, that as the derivative order increases, the risk of precision loss grows. Despite depending on the specific function being differentiated, our numerical experiments show that derivatives of orders around 100 may require quadruple precision for accurate computation. For even higher-order derivatives, further increases in precision will likely be necessary.
Our Fortran implementation uses complex numbers as the components of dual numbers. Although an implementation with real numbers may be necessary for efficiency reasons, we chose to use complex numbers for greater generality. It is important to note the lack of a consistent convention regarding the branches chosen for complex functions across different programming environments. For example, when computing $\arcsin(1.1)$, the result obtained from SBCL (version 2.2.9.debian), Octave (version 7.3.0), and Wolfram Mathematica (version 12.0) is $(1.5708,-0.44357)$. In contrast, the same computation using Python’s cmath library yields $(1.5708,0.44357)$, a result consistent with that produced by modern versions of Fortran. We have elected to adopt the latter convention, aligning with the Fortran implementation, to avoid the necessity of developing custom functions to select a specific branch. The decision regarding which branch to standardize will ultimately rest with the Fortran standards committee (ISO/IEC JTC1/SC22/WG5).
2 Dual numbers and derivatives
2.1 First order case
Analogous to the definition of a complex number $z=a+i\,b$ where $a$, $b$ are real numbers and $i$, the imaginary unit satisfying $i^{2}=-1$, a dual numbers is defined as
$r$ $=a_{0}\,\epsilon_{0}+\epsilon_{1}\,a_{1},$ (1)
$r$ $=a_{0}+\epsilon_{1}\,a_{1},$ (2)
where $a_{0}$, $a_{1}$ are real or complex numbers, $\epsilon_{0}=1$, and $\epsilon_{1}$ is the dual unit satisfying
$\epsilon_{1}^{2}=0.$ (4)
In a similar manner to how an analytical real function is extended to a complex function, we can evaluate the Taylor expansion of a function $f(z)$ at $z+\epsilon_{1}$:
$f(z+\epsilon_{1})=f(z)+f^{\prime}(z)\epsilon_{1}+\frac{1}{2}f^{\prime\prime}(z)\epsilon_{1}^{2}+\cdots.$ (5)
However, in virtue of Eq. (4) $\epsilon_{1}^{k}=0$ for $k>1$. This simplifies the expansion to
$f(z+\epsilon_{1})=f(z)+f^{\prime}(z)\epsilon_{1}.$ (6)
Therefore, evaluating an analytic function at the dual number $z+\epsilon_{1}$ yields a dual number where its $\epsilon_{0}$ component is $f(z)$ and its $\epsilon_{1}$ component is $f^{\prime}(z)$. The process of extending $f(z)$ to operate on dual numbers is called dualizing $f(z)$. As a simple example, the dual extension of the sine function $sin(z)$ will be
$sin(z)=\sin z+\cos z\;\epsilon_{1}.$ (7)
Here, the function name in italics indicates the dual version of the original function. It is important to note that this function is not the general dual version for the “$\sin$” function, as $z$ is a complex number. To construct the general dual version we need to be able of computing $sin(g)$ for $g=g_{0}+\epsilon_{1}\,g_{1}$ a dual number. Setting $g(z+\epsilon_{1})$ = $g_{0}+g_{1}\epsilon_{1}$ and evaluating in the Taylor series expansion for the function $f$, we have
$f(g_{0}+\epsilon_{1}\,g_{1})=f(g_{0})+f^{\prime}(g_{0})g_{1}\,\epsilon_{1}.$ (8)
Therefore, the dual extension of the sine function becomes:
$sin(g)=\sin(g_{0})+\cos(g_{0})\,g_{1}\,\epsilon_{1}.$ (9)
This procedure can be applied to dualize all elementary functions [^1] A non-elementary function, such as the error function $\text{erf}(x)$, can be extended to dual numbers in the same way. However, since this function is not implemented in Fortran for complex values, it is not extended in our dual number implementation. and main operators of a programming language.
The generalization for calculating arbitrary-order derivatives is straightforward from a theoretical standpoint. However, computational implementation presents significant challenges, particularly in avoiding stack overflow errors that can occur at higher derivative orders when using recursive or nested structures. Most works in the literature focus on implementing dual numbers for the real case, typically up to the first or second order [20, 21, 22, 23, 19]. Some extensions to third and fourth orders, including the complex case, have also been proposed [16, 17]. While implementations using recursive or nested structures are effective and efficient for lower-order derivatives, they may encounter limitations at very high derivative orders due to stack depth limitations. Achieving dual number implementations for arbitrary orders without relying on recursion is the primary contribution of our study.
2.2 arbitrary order case
Derivatives to arbitrary order can be calculated by defining the dual number [17]
$r_{n}=\sum_{k=0}^{n}a_{k}\epsilon_{k},$ (10)
with $a_{k}$ a complex number and
$\epsilon_{i}\cdot\epsilon_{j}=\begin{cases}0&\mbox{if }i+j>n,\vspace*{0.3cm}\\ \frac{(i+j)!}{i!\,j!}\epsilon_{i+j}&\mbox{otherwise. }\end{cases}$ (11)
Now, evaluating the Taylor expansion for an analytical function in the number $z+\epsilon_{1}$ and using Eq.(11) we have
$f(z+\epsilon_{1})=f(z)\epsilon_{0}+f^{\prime}(z)\epsilon_{1}+\dots+f^{(n)}(z)\epsilon_{n}.$ (12)
Computing Eq. (12) for the elementary function is not very complicated, and thus the dual extension for $f(z)$ is relatively easy. However, in the general case, the dual version for $f(g(z))$, i.e. $f(g)$ is much more complicated. Nevertheless, the process is simplified if one first implements the well known Faà Di Bruno formula [24, 25] for dual numbers as it is showed in the next section.
3 Dual number implementation to arbitrary order
To implement dual numbers to arbitrary order, we will first implement the chain rule to calculate the $n$-th derivative of a function in Fortran. The data structure we will use is shown in Listing 1.
1 type, public :: dualzn
2 complex(prec), allocatable, dimension(:) :: f
3 end type dualzn
Listing 1: Fortran Code for dualzn Type using the precision given by prec. The $k$-th component of a dual number $g$ of this type is accessed by $g\%f(k)$.
Now, according to the famous Faà Di Bruno formula, the $n$-th derivative of $f(g(x))$ is given by
$D^{n}f(g(x))=\sum\frac{n!}{k_{1}!k_{2}!\cdots k_{n}!}f^{(k_{1}+k_{2}+\cdots+k_{n})}(g(x))\prod_{j=1}^{n}\left(\frac{g^{(j)}(x)}{j!}\right)^{k_{j}},$ (13)
where $D^{n}=d^{n}/dx^{n}$, $f^{(k)}(x)=D^{k}f(x)$ and the the sum is taken over all sequences of nonnegative integers $k_{1},k_{2},\ldots,k_{n}$ such that they obey the Diophantine equation:
$k_{1}+2k_{2}+3k_{3}+\cdots+nk_{n}=n.$ (14)
Although a recursive formula that avoid solving the Diophantine equation can be obtained [26], a more direct approach is to write Eq. (13) as:
$D^{n}f(g(x))=\sum_{k=1}^{n}f^{(k)}(g(x))\;B_{n,k}(g^{\prime}(x),g^{\prime\prime}(x),\dots,g^{(n-k+1)}(x)),$ (15)
where the $B_{n,k}(x_{1},x_{2},\dots,x_{n-k+1})$ are the partial Bell polynomials which can be recursively defined by [27]
$B_{n,k}(x_{1},x_{2},\dots,x_{n-k+1})=\sum_{i=0}^{n-k}\binom{n-1}{i}x_{i+1}B_{n-i-1,k-1}(x_{1},x_{2},\dots,x_{n-k-i+1}),$ (16)
with
$B_{0,0}(\cdot)$ $=1$ (17)
$B_{n,0}(\cdot)$ $=0\text{~{}for~{}}n\geqslant 1$ (18)
$B_{0,k}(\cdot)$ $=0\text{~{}for~{}}k\geqslant 1.$ (19)
From this recursive definition, we can build the iterative version shown in Algorithm 3.
[…]
1:$n,k\in\mathbb{Z}$, $x\in\mathbb{C}^{m}$
2:$result\_value\in\mathbb{C}$
3:Initialize $dp[n+1,k+1]\leftarrow 0$
4:$dp[1,1]\leftarrow 1$
5:for $nn\leftarrow 1$ to $n$ do
6:     $dp[nn+1,1]\leftarrow 0$
7:end for
8:for $kk\leftarrow 1$ to $k$ do
9:     $dp[1,kk+1]\leftarrow 0$
10:end for
11:if $n=0$ and $k=0$ then
12:     $result\_value\leftarrow 1$
13:     return $result\_value$
14:else if $\text{size}(x)=0$ then
15:     $result\_value\leftarrow 0$
16:     return $result\_value$
17:end if
18:for $nn\leftarrow 1$ to $n$ do
19:     for $kk\leftarrow 1$ to $k$ do
20:         $LX\leftarrow\text{size}(x)$
21:         $nx\leftarrow\max(nn-kk+1,LX)$
22:         if $nx>0$ then
23:              Allocate $newx[nx]\leftarrow 0$
24:              $newx[1:LX]\leftarrow x$
25:              $sum\_val\leftarrow 0$
26:              for $ii\leftarrow 0$ to $nn-kk$ do
27:                  $sum\_val\leftarrow sum\_val+\binom{nn-1}{ii}\cdot newx[ii+1]\cdot dp[nn-ii,kk]$
28:              end for
29:              $dp[nn+1,kk+1]\leftarrow sum\_val$
30:              Deallocate $newx$
31:         end if
32:     end for
33:end for
34:$result\_value\leftarrow dp[n+1,k+1]$
35:return $result\_value$
Having implemented the $B_{n,k}$ function, we can now implement the chain rule Eq. (15) in Fortran using the derived data type dualzn from Listing 1 and following Algorithm 3. Here, the $fci$ function provides the dual version of $f(z)$. We emphasize that it is not the dual version of $f(g(z))$, which is constructed after applying the chain rule to $f$ through function Dnd. Let us exemplify this with the sinusoidal function.
[…]
1:$fci$: Function, $gdual$: dualzn type, $n\in\mathbb{Z}$
2:$dnfcg\in\mathbb{C}$
3:$g0\leftarrow gdual\%f(0)$
4:$fvd\leftarrow fci(g0)$
5:if $n=0$ then
6:     $dnfcg\leftarrow fvd\%f(0)$
7:else
8:     $suma\leftarrow 0$
9:     for $k\leftarrow 1$ to $n$ do
10:         Allocate $xvg[1:n-k+1]$
11:         for $j\leftarrow 1$ to $n-k+1$ do
12:              $xvg[j]\leftarrow gdual\%f(j)$
13:         end for
14:         $suma\leftarrow suma+fvd\%f(k)\cdot B(n,k,xvg)$
15:         Deallocate $xvg$
16:     end for
17:     $dnfcg\leftarrow suma$
18:end if
19:return $dnfcg$
The dual version $sin(z)$ to order $n$, of the $\sin(z)$ function is given by
$sin(z)$ $=\sum_{k=0}^{n}D^{k}(\sin z)\,\epsilon_{k}$
$sin(z)$ $=\sum_{k=0}^{n}\sin\left(z+k\,\frac{\pi}{2}\right)\epsilon_{k}.$ (20)
We recall that in this function $sin\;z$ is a dual number but $z$ is not. To construct the dual version $sin(g)$, being $g$ a dual number, we need to apply the Dnd function to $sin\;z$. Therefore
$sin(g)=\sum_{k=0}^{n}\text{Dnd}(sin,g,k)\,\epsilon_{k}$ (21)
will be the dual version of $\sin(g(z))$.
By following this procedure, the dual version for all elementary functions can be obtained. As exemplified in Eq. (20), to obtain the dual version for $f(z)$ its derivatives are necessary. For most of the elementary functions this can be done with no much effort. There are another functions, however, for which to obtain its dual version $f(z)$ could be a little more complicated. Take for instance the inverse of the sine function, the $\arcsin(z)$ function. A recursive [^2] While a recursive approach to calculating the derivatives of functions may appear more direct, we advise against it due to potential inefficiencies, including lower performance and higher memory consumption. formula to calculate its $n$-th derivative is given by
$D^{0}(\arcsin z)$ $=\arcsin z$ (22)
$D^{1}(\arcsin z)$ $=\frac{1}{\sqrt{1-z^{2}}}$ (23)
$D^{n}(\arcsin z)$ $=-\sum_{k=1}^{n-1}\binom{n-1}{k}D^{k}\left(\sqrt{1-z^{2}}\right)\frac{D^{n-k}(\arcsin z)}{\sqrt{1-z^{2}}}.$ (24)
One can pursue this approach; however, if we first dualize the functions $inv(z)=1/z$ as well as the square root function and the difference and multiplication operators, then by expressing Eq. (23) in its dual form, all higher-order derivatives can be automatically computed. All these function are relatively easy to dualize, for instance, Algorithms 3 and 4 shows the dualization of the multiplication operator using the Leibniz rule.
Algorithm 3 Pseudocode for the function timesdzn(A,B,k), the Leibniz rule.
1:$A$: dualzn type, $B$: dualzn type, $k\in\mathbb{Z}$
2:$fr\in\mathbb{C}$
3:$fr\leftarrow 0$
4:for $i\leftarrow 0$ to $k$ do
5:     $fr\leftarrow fr+\binom{k}{i}\cdot A\%f(i)\cdot B\%f(k-i)$
6:end for
7:return $fr$
Algorithm 4 Pseudocode for the function timesd(A,B) which implements the product of the dual numbers A and B
1:$A$: dualzn type, $B$: dualzn type
2:$fr$: dualzn type
3:Allocate $fr\%f(0:\text{order})$
4:for $k\leftarrow 0$ to order do
5:     $fr\%f(k)\leftarrow\texttt{timesdzn}(A,B,k)$
6:end for
7:return $fr$
4 The DNAOAD package
With the just presented theory, dual numbers to arbitrary order are implemented in the so-called DNAOAD package. This implementation not only allows for computation of derivatives also to arbitrary order, but also allows the construction of algorithms in the context of dual numbers. Before presenting some interesting applications we first describe the DNAOAD package.
4.1 Elements of the package
The implementation is available for double (real64) and quadruple (real128) precisions. The user can choose which precision to use. The module for double precision is shown in Listing 2. This module is used by the dualzn_mod module, which is the core of the dual number implementation and contains approximately 2,000 lines of code. The first 28 lines, which also show the functions and operators extended to dual numbers (overloaded), are shown in Listing 3.
1module precision_mod
2 use iso_fortran_env, only: int32, int64, real32, real64, real128
3 implicit none
4
5 integer, parameter :: prec = real64
6end module precision_mod
Listing 2: The precision module for double precision implementation.
1module dualzn_mod
2 use precision_mod
3 implicit none
4
5 private
6 !---------------------------------------------------------------------
7 !Some Module variables
8 !Default order, can be modified with set_order
9 integer, public :: order = 1
10 real(prec), public, parameter :: Pi = 4.0_prec*atan(1.0_prec)
11 !---------------------------------------------------------------------
12
13 !dual number definition to any order
14 type, public :: dualzn
15 complex(prec), allocatable, dimension(:) :: f
16 end type dualzn
17
18 public :: set_order, initialize_dualzn, f_part
19 public :: binomial, BellY, Dnd
20 public :: itodn, realtodn, cmplxtodn, Mset_fpart
21
22 public :: inv, sin, cos, tan, exp, log, sqrt, asin, acos, atan, asinh
23 public :: acosh, atanh, sinh, cosh, tanh, absx, atan2
24 public :: conjg, sum, product, matmul
25
26 public :: assignment (=)
27 public :: operator(*), operator(/), operator(+), operator(-)
28 public :: operator(**), operator(==), operator(/=)
Listing 3: Extract from the module dualzn_mod
The description of the variables, functions and operators in listing 3 is as follows.
  1. 1.
    order: this define a module variable which define the order of the dual number to work with. The default is 1 but can be changed to any desired integer greater than 0 using the set_order(n) subroutine explained below.
  2. 2.
    set_order(n): Subroutine that sets the order of the dual numbers.
    Argument:
    • n, an integer number.
  3. 3.
    initialize_dualzn(r): elemental (element-wise operations) subroutine which initialize to zero a dual quantity.
    Argument:
    • r, a type(dualzn) quantity than can be scalar, or array.
    Output: a dual quantity, scalar or array.
  4. 4.
    f_part(r,k): elemental function which extract the k-th dual part of the dual quantity r.
    Arguments:
    • r, a type(dualzn) quantity than can be scalar, or array.
    • k, an integer number.
    Output: a dual quantity, scalar or array.
    This Function if for the user convenience as the s%f(k) operation can also be used to access the k-th part of a the scalar dual number s.
  5. 5.
    binomial(m,n): returns the binomial coefficient, representing the number of ways to choose n elements from a set of m elements.
    Arguments:
    • m, n, integers.
    Output: a real number, even when by definition $\binom{m}{n}$ in an integer number.
  6. 6.
    BellY(n, k, z): function that computes the partial Bell polynomials $B_{n,k}$.
    Arguments:
    • n, k, integers.
    • z, an array of complex numbers, the point of evaluation.
    Output: a complex number.
  7. 7.
    Dnd(fc,gdual,n): function that implements the Faà Di Bruno’s formula, the chain rule to calculate $D^{n}(f(g(x)))$.
    Arguments:
    • fc, a function of type procedure(funzdual) with funzdual given in the abstract interface:
      1abstract interface
      2 pure function funzdual(z_val) result(f_result)
      3 use precision_mod
      4 import :: dualzn
      5 complex(prec), intent(in) :: z_val
      6 type(dualzn) :: f_result
      7 end function funzdual
      8end interface
    • gdual, a type(dualzn) number.
    • n, an integer.
    Output: a complex number.
    Although essential to the dual number implementation, this function is not meant for regular use, except if the user wants to dualize their own functions.
  8. 8.
    itodn(i), realtodn(x), cmplxtodn(z): functions that promote an integer, real, and complex number to a dualzn number. Since the assignment operator (=) is overloaded, these functions may be of less use.
  9. 9.
    Mset_fpart(k,cm,A): this subroutine sets the dual-k component of matrix A to cm
    Arguments:
    • k, integer.
    • cm, a complex number.
    • A, a dualzn matrix.
    Output: the matrix A.
  10. 10.
    sin, cos, tan, exp,  log,  sqrt,  asin,  acos,  atan, asinh, acosh, atanh, sinh,
    cosh, tanh, atan2, conjg, sum, product, matmul, =, *, /, +, -, **, ==, /=
    : are the same Fortran functions and operators overloaded to deal with arguments of the type dualzn numbers.
  11. 11.
    inv(r): the inverse of a dualzn number, equivalent to 1/r.
    Arguments:
    • r, a type(dualzn) number.
    Output: a type(dualzn) number.
  12. 12.
    absx(r): absx(r) = sqrt(r*r) is not sqrt(r*conjg(r)). This function is coded to be used (if necessary) with the complex steep approximation method [28, 29], is not the overloaded abs function, except for the real case.
    Arguments:
    • r, a type(dualzn) number.
    Output: a type(dualzn) number.
Additionally to the already discussed modules (precision_mod and dualzn_mod) the package also contains the module diff_mod which contains some useful differntial operators. The components (interfaces and functions) of this module are described below.
  1. 1.
    fsdual: Abstract interface for a scalar dual function $f:\mathbb{D}^{m}\to\mathbb{D}$ defined by
    abstract interface
       function fsdual(xd) result(frsd)
         use dualzn_mod
         type(dualzn), intent(in), dimension(:) :: xd
         type(dualzn) :: frsd
       end function fsdual
    end interface
  2. 2.
    fvecdual: Abstract interface for a vector dual function $f:\mathbb{D}^{m}\to\mathbb{D}^{n}$ defined by
    abstract interface
       function fvecdual(xd) result(frd)
         use dualzn_mod
         type(dualzn), intent(in), dimension(:)  :: xd
         type(dualzn), allocatable, dimension(:) :: frd
       end function fvecdual
    end interface
  3. 3.
    dfv = d1fscalar(fsd,v,q)
    dfv: complex(prec). First-order directional derivative of a scalar function along vector v, evaluated at point q.
    fsd: procedure(fsdual). Is a scalar
    dualzn function $f:\mathbb{D}^{m}\to\mathbb{D}$ (similar to $f:\mathbb{R}^{m}\to\mathbb{R}$).
    v: complex(prec), intent(in), dimension(:). Vector along which the directional derivative will be computed.
  4. 4.
    d2fv = d2fscalar(fsd,v,q)
    d2fv: complex(prec). Second-order directional derivative of the scalar function $f:\mathbb{D}^{m}\to\mathbb{D}$, along vector v, evaluated at point q.
    fsd: procedure(fsdual). Is a scalar
    dualzn function $f:\mathbb{D}^{m}\to\mathbb{D}$.
    v: complex(prec), intent(in), dimension(:). Vector along which the directional derivative will be computed.
    q: complex(prec), intent(in), dimension(:). Is the evaluating point.
    Note: This function is equivalent to $\mathbf{v.H.v}$, the product of the Hessian matrix with vector $\mathbf{v}$, but with higher efficiency.
  5. 5.
    d2fv = d2fscalar(fsd,u,v,q)
    d2fv: complex(prec). Second-order directional derivative of the scalar function $f:\mathbb{D}^{m}\to\mathbb{D}$, along vectors u, v, evaluated at point q.
    fsd: procedure(fsdual). Is a scalar
    dualzn function $f:\mathbb{D}^{m}\to\mathbb{D}$.
    u, v: complex(prec), intent(in), dimension(:). Vectors along which the directional derivative will be computed.
    q: complex(prec), intent(in), dimension(:). Is the evaluating point.
    Note: This function is equivalent to $\mathbf{u.H.v}$, the product of the Hessian matrix with vectors $\mathbf{u}$ and $\mathbf{v}$, but with higher efficiency.
  6. 6.
    dfvecv = d1fvector(fvecd,v,q,n)
    dfvecv: complex(prec), dimension(n). Second-order directional derivative of the vector function $f:\mathbb{D}^{m}\to\mathbb{D}^{n}$, along vector v, evaluated at point q. fvecd: procedure(fvecdual). Is a vector dualzn function $f:\mathbb{D}^{m}\to\mathbb{D}^{n}$.
    v: complex(prec), intent(in), dimension(:). Vector along which the directional derivative will be computed.
    q: complex(prec), intent(in), dimension(:). Is the evaluating point.
    Note: This function is equivalent to $\mathbf{J.v}$, the product of the Jacobian matrix with vector $\mathbf{v}$, but with higher efficiency.
  7. 7.
    H = Hessian(fsd,q)
    H: complex(prec), dimension (size(q),size(q)). The hessian matrix.
    fsd: procedure(fsdual). Is a scalar
    dualzn function $f:\mathbb{D}^{m}\to\mathbb{D}$.
    q: complex(prec), intent(in), dimension(:). Is the evaluating point.
  8. 8.
    J = Jacobian(fvecd,q,n)
    J: complex(prec), dimension(n,size(q)). The Jacobian matrix.
    fvecd: procedure(fvecdual). Is a vector dual function $f:\mathbb{D}^{m}\to\mathbb{D}^{n}$.
    q: complex(prec), intent(in), dimension(:). Is the evaluating point.
    n: integer, intent(in). The dimension of
    fvecd.
  9. 9.
    G = gradient(fsd,q)
    G: complex(prec), dimension(size(q)). The gradien vector.
    fsd: procedure(fsdual). Is a scalar
    dualzn function $f:\mathbb{D}^{m}\to\mathbb{D}$.
    q: complex(prec), intent(in), dimension(:). Is the evaluating point.
4.2 Usage of DNAOAD
The package is available for Windows and GNU/Linux in both double and quadruple precision, and it supports the gfortran and ifx Intel Fortran compilers. It can be downloaded from [30]. After uncompressing the file, the user can navigate, for instance, to the folder:
*/DNAOAD/GNU_Linux/ifx_case/prec_real64/examples/’ (the ‘*’ depends on where the package is uncompressed on the local machine), and run some of the examples. The source code is also provided; however, in our examples, we will assume that the precompiled libraries in the GNU_Linux folder are used.
4.2.1 Simple examples
As a simple example, the program in listing 4 computes the derivatives from the 0-th up to the 5-th order of $f=\sin(z)^{\log(z^{2})}$ at $z_{0}=1.1+2.2i$. Assuming the code is saved in the folder ‘*/DNAOAD/GNU_Linux/ifx_case/prec_real64/examples/’ under the file example1.f90, it can be compiled in GNU/Linux as:
ifx -I../LibDualzn64 -o example1.out example1.f90 -L../LibDualzn64 -ldualzn
and in Windows as:
ifx -o example1.exe example1.f90 -I../LibDualzn64 ../LibDualzn64/libdualzn.lib
For the gfortran compiler, use:
gfortran -I../LibDualzn64 -o example1.exe example1.f90 -L../LibDualzn64 -ldualzn
Alternatively, run the provided script create_exec.sh (on GNU/Linux) or create_exec.bat (on Windows), followed by the file name where the code is saved.
1program main
2 use precision_mod
3 use dualzn_mod
4 implicit none
5 complex(prec) :: z0
6 type(dualzn) :: r, fval
7 integer :: k
8
9 call set_order(5) !we set the order to work with
10 r = 0 !we initialize the dual number to 0, alternativelly:
11 !call initialize_dualzn(r)
12
13 z0 = (1.1_prec,2.2_prec) !the evaluating point
14
15 !we set the 0-th and 1-th components. If dual numbers are used to
16 !calculate D^n f(z0) then r must be of the form r = r0 + 1*eps_1
17 r%f(0) = z0
18 r%f(1) = 1
19 fval = sin(r)**log(r*r)
20 !writing the derivatives, from the 0th derivative up to the
21 !order-th derivative.
22 print*,"derivatives"
23 do k=0,order
24 print*,fval%f(k)
25 end do
26end program main
Listing 4: Example of derivative calculation. Since a dualzn number is an allocatable entity, it must first be initialized.
Dual numbers can also be used to differentiate Fortran code. Consider the function $f(x)=\sin(x)\exp(-x^{2})$ and the task of calculating the derivatives of $g=f(f(\cdots f(x)\cdots))$, where $f$ is nested 1,000 times. This calculation is virtually impossible symbolically, and finite differences would be inefficient and inaccurate. Listing 5 shows a program to compute these derivatives. Appendix A., presents Python and Julia versions when the funcion is nested 5 times and the order of derivation is 10. While this example is theoretical, real-world scenarios often involve multiple function compositions, vector rotations, and similar operations. For a practical demonstration, refer to [16], where kinematic quantities for the coupler point in a spherical 4R mechanism are calculated.
A notable aspect of this illustration is the computation of the 15th-order derivative. Although physical problems rarely require derivatives beyond the fourth order, the ability to compute higher-order derivatives remains valuable due to potential future applications. Therefore, the importance of being able to handle such calculations should not be dismissed.
1program main
2 use precision_mod
3 use dualzn_mod
4 implicit none
5
6 complex(prec) :: z0
7 type(dualzn) :: r, fval
8 integer :: k
9 real :: t1,t2
10
11 call set_order(15) !we set the order to work with
12
13 !since a dualzn numbers is an allocatable entity, do not forget to
14 !initialize it
15 r = 0 !<--- initializing r to 0
16 r%f(0) = (1.1_prec,0.0_prec)
17 r%f(1) = 1 !since we want to differentiate, r = r0 +1*eps_1
18 !all the other components are 0 as r was initialized to 0
19
20 call cpu_time(t1)
21 fval = ftest(r)
22 call cpu_time(t2)
23
24 !Writing the derivatives, from the 0th derivative up to the
25 !order-th derivative.
26 print*,"derivatives"
27 do k=0,order
28 write(*,"(i0,a,f0.1,a,e17.10)"),k,"-th derivative at x = ", &
29 real(r%f(0)),":",real(fval%f(k))
30 end do
31
32 print*,"elapsed time (s):",t2-t1
33
34contains
35 function ftest(x) result(fr)
36 type(dualzn), intent(in) :: x
37 type(dualzn) :: fr
38 integer :: k
39
40 !nested function f(x) = sin(x) * exp(-x^2) f(f(...(f(x))...))
41 !applied 1000 times
42 fr = sin(x)*exp(-x*x)
43 do k=1, 1000-1
44 fr = sin(fr)*exp(-fr*fr)
45 end do
46 end function ftest
47end program main
Listing 5: Example of differentiating a function that is not given in closed form but implemented as a computer program.
4.2.2 Calculating gradients, Jacobians and Hessians
The package also includes the diff_mod module, which provides useful functions for computing gradients, Jacobians, and Hessians. The underlying theory of using dual numbers to compute these differential operators, as well as directional derivatives in general, is presented in [17]. Below an example of use of this module. This example can be found in the folder:
*/DNAOAD/GNU_Linux/ifx_case/prec_real64/examples/’.
1!module with example of functions
2module function_mod
3 use dualzn_mod
4 implicit none
5 private
6
7 public :: fstest, fvectest
8
9contains
10 !Example of scalar function f = sin(x*y*z) + cos(x*y*z)
11 function fstest(r) result(fr)
12 type(dualzn), intent(in), dimension(:) :: r
13 type(dualzn) :: fr
14 type(dualzn) :: x,y,z
15
16 x = r(1); y = r(2); z = r(3)
17 fr = sin(x*y*z) + cos(x*y*z)
18 end function fstest
19
20 !Example of vector function f = [f1,f2,f3]
21 !f = fvectest(r) is a function f:D^m ---> Dn
22 function fvectest(r) result(fr)
23 type(dualzn), intent(in), dimension(:) :: r
24 type(dualzn), allocatable, dimension(:) :: fr
25 type(dualzn) :: f1,f2,f3
26 type(dualzn) :: x,y,z,w
27
28 x = r(1); y = r(2); z = r(3); w = r(4)
29
30 f1 = sin(x*y*z*w)
31 f2 = cos(x*y*z*w)*sqrt(w/y - x/z)
32 f3 = sin(log(x*y*z*w))
33
34 allocate(fr(3))
35 fr = [f1,f2,f3]
36 end function fvectest
37end module function_mod
38
39!main program
40program main
41 use precision_mod
42 use dualzn_mod
43 use diff_mod
44 use function_mod
45 implicit none
46
47 integer, parameter :: nf =3, mq = 4
48 complex(prec), parameter :: ii = (0,1)
49 complex(prec), dimension(mq) :: q, vec
50 complex(prec), dimension(nf,mq) :: Jmat
51 complex(prec), dimension(nf) :: JV
52 complex(prec), dimension(3) :: GV
53 complex(prec), dimension(3,3) :: Hmat
54 integer :: i
55
56 vec = [1.0_prec,2.0_prec,3.0_prec,4.0_prec]
57 q = vec/10.0_prec + ii
58
59 print*,"Jv using matmul"
60 Jmat = Jacobian(fvectest, q , nf)
61 JV = matmul(Jmat,vec)
62 do i=1,nf
63 write(*,*) JV(i)
64 end do
65 write(*,*)
66
67 print*,"Jv using vector directional derivative"
68 JV = d1fvector(fvectest,vec,q,nf)
69 do i=1,nf
70 write(*,*) JV(i)
71 end do
72 write(*,*)
73
74 print*,"---Hessian matrix---"
75 Hmat = Hessian(fstest,q(1:3))
76 do i=1,3
77 write(*,"(A,i0)") "row:",i
78 write(*,*) Hmat(i,:)
79 end do
80 write(*,*)
81
82 print*,"---Gradient---"
83 GV = gradient(fstest,q(1:3))
84 do i=1,3
85 write(*,*) GV(i)
86 end do
87end program main
Listing 6: Example of using the diff_mod.
5 Conclusions
This study addresses the challenges of computing higher-order derivatives with dual numbers by introducing DNAOAD, a Fortran-based implementation designed for arbitrary-order automatic differentiation. Most existing dual-number implementations are limited to lower-order derivatives, with only one notable approach—Julia’s ForwardDiff—supporting arbitrary-order differentiation through nested dual numbers. However, no prior work has demonstrated a method for arbitrary-order differentiation with dual numbers that avoids reliance on nested structures, which are prone to significant memory consumption and stack overflow errors. DNAOAD bridges this gap by directly implementing dual numbers to arbitrary orders, eliminating the need for recursive or nested frameworks. This direct and innovative approach significantly reduces computational overhead, enabling the efficient and reliable computation of derivatives at very high orders, and providing a scalable and practical solution to the limitations of existing methods.
Through numerical experiments, DNAOAD demonstrated its capability to compute derivatives at very high orders with remarkable accuracy, though extremely high orders may require enhanced numerical precision, such as quadruple precision. The implementation also provides a considerable number of useful functions for various applications, including the calculation of gradients, Jacobians, and Hessians, as well as higher-order derivatives of complex functions. These capabilities enhance DNAOAD’s applicability to a broad range of fields, such as optimization, physics, and engineering, where precise differentiation is essential.
In comparison with existing tools, such as PyTorch and ForwardDiff—which excel in low-order differentiation but face scalability issues due to memory overheads—DNAOAD provides a scalable and efficient alternative. Its robustness was demonstrated in experiments involving deeply nested functions, highlighting its potential for tackling complex computational tasks. This implementation represents a significant advancement in practical automatic differentiation by combining theoretical rigor with computational efficiency. Future work may focus on extending DNAOAD to other programming environments and further optimizing its performance for large-scale applications.
Data availability statement
The Fortran source code, which contains the data supporting the findings of this study, is openly available on GitHub at https://github.com/fpenunuri/DNAOAD.
Appendix Appendix A. Computing higher order derivatives with PyTorch and ForwardDiff
Although the computation of higher-order derivatives can lead to significant memory consumption, the utility of the PyTorch and ForwardDiff libraries remains undeniable. Below, we present the example of Listing 5 that demonstrates a relatively small number of function compositions in both Python and Julia, respectively.
1import torch
2import time
3
4#nested function f(x) = sin(x) * exp(-x^2) applied 5 times
5def ftest(x):
6 fr = torch.sin(x) * torch.exp(-x**2)
7 for _ in range(5 - 1):
8 fr = torch.sin(fr)*torch.exp(-fr**2)
9 return fr
10
11#value for x
12x = torch.tensor(1.1, dtype=torch.float64, requires_grad=True)
13
14start = time.time()
15#computing the nested function
16grad_0 = ftest(x)
17
18order = 10 #order > 10 probably led to a crash.
19#Computing derivatives
20grad_k=[grad_0]
21for k in range (order+1):
22 grad_k.append(torch.autograd.grad(grad_k[k], x, create_graph=True)[0])
23 print(f"{k}th-order derivative at x = {x.item()}: {grad_k[k].item()}")
24
25end = time.time()
26
27print("Elapsed time (s):", end - start)
Listing 7: Example of computing higher-order derivatives with PyTorch.
1using ForwardDiff
2using BenchmarkTools
3
4#The function
5function ftest(x)
6 fr = sin(x) * exp(-x^2)
7 for _ in 1:5-1
8 fr = sin(fr) * exp(-fr^2)
9 end
10 return fr
11end
12
13#Function to compute nth derivative using ForwardDiff and Float64
14function nth_derivative(f, x::Float64, n::Int)
15 if n == 0
16 return f(x)
17 elseif n == 1
18 return ForwardDiff.derivative(f, x)
19 else
20 return nth_derivative(y -> ForwardDiff.derivative(f, y), x, n - 1)
21 end
22end
23
24#input value
25x = 1.1
26
27#println("")
28order = 10 #order > 14 probably led to a crash.
29elapsed_time = @elapsed begin
30 for n in 0:order
31 derivative_n = nth_derivative(ftest, x, n)
32 println("\$n-th derivative at x = \$x: \$derivative_n")
33 end
34end
35
36println("Elapsed time: \$elapsed_time seconds")
Listing 8: Example of computing higher-order derivatives with ForwardDiff.

### v2

@@
 Abstract
-Dual numbers are a well-known tool for computing derivatives of functions.
-While the theoretical framework for calculating derivatives of arbitrary order is well established, practical implementations remain less developed.
-One notable implementation is available in the Julia programming language where dual numbers are designed to be nested, enabling the computation of derivatives to arbitrary order.
-However, this approach has a significant drawback as it struggles with scalability for high-order derivatives.
-The nested structure quickly consumes memory, making it challenging to compute derivatives of higher orders.
-In this study, we introduce DNAOAD, a Fortran-based implementation of automatic differentiation capable of handling derivatives of arbitrary order using dual numbers.
-This implementation employs a direct approach to represent dual numbers without relying on recursive or nested structures.
-As a result, DNAOAD facilitates the efficient computation of derivatives of very high orders while addressing the memory limitations of existing methods.
+Dual numbers are a well-established tool for computing derivatives and constitute the basis of forward-mode automatic differentiation.
+While the theoretical framework for computing derivatives of arbitrary order is well understood, practical and scalable implementations remain limited.
+Existing approaches based on nested dual numbers, such as those used in modern high-level languages, suffer from severe memory growth and poor scalability as the derivative order increases.
+In this work, we introduce DNAOAD, a Fortran-based automatic differentiation framework capable of computing derivatives of arbitrary order using dual numbers with a direct, non-nested representation.
+By avoiding recursive data structures, DNAOAD significantly reduces memory usage and enables the efficient computation of derivatives of very high order, overcoming key scalability limitations of existing methods and making it particularly well suited for high-performance scientific computing applications.
 Facultad de Ingeniería, Universidad Autónoma de Yucatán, A.P. 150, Cordemex, Mérida, Yucatán, México.
@@
 1 Introduction
-Differentiation is a fundamental process that is widely applied in numerous fields of science and engineering.
-It can be carried out symbolically using specialized software, such as Maxima, Maple, or Mathematica, or numerically through finite-difference methods.
-Another approach is Automatic Differentiation (AD) [1, 2, 3], which enables the numerical computation of derivatives with the same level of accuracy and precision as symbolic methods, but without requiring symbolic computations.
-Automatic Differentiation (AD) is a systematic algorithmic technique for efficiently applying the chain rule to compute derivatives.
-Although there is extensive research on this topic, most studies focus on the computation of first- and second-order derivatives for real-valued functions [4, 5, 6, 7, 8, 9, 10].
-ADOL-C [11, 12, 13] is an interesting implementation of this technique, which takes advantage of operator overloading to enable differentiation.
-Although ADOL-C supports higher-order derivatives for some functions, a considerable number, including asin, acos, atan, and exponentiation, lack full support for differentiation across all derivative orders.
-Notable AD implementations that support the computation of derivatives of arbitrary order include the PyTorch library in Python [14] and the ForwardDiff package in Julia [15].
-Even when these libraries do not natively support automatic differentiation for complex-valued functions, the complex case can be managed by separately computing the derivatives of the real and imaginary components of the functions involved.
-PyTorch is a versatile and widely used Python library for AD and deep learning applications.
-One of its key strengths is its ability to compute higher-order derivatives using a dynamic computation graph.
-This graph tracks operations at runtime and facilitates automatic computation of gradients.
-However, when dealing with higher-order differentiation—especially when calculating derivatives of complex operations or multiple gradients—PyTorch retains the entire computational graph in memory.
-While this is necessary to compute successive gradients, it significantly increases memory consumption.
-For instance, calculating higher-order derivatives can cause the computational graph to rapidly expand due to the accumulation of intermediate values.
-Consequently, PyTorch’s memory usage may quickly exceed the system’s available memory, often resulting in a std::bad_alloc error, which occurs when memory allocation requests surpass the system’s capacity.
-In our numerical experiment, performed on a 12th generation Intel(R) Core(TM) i7-12650H CPU @ 4.70 GHz (base 400 MHz) with 15 GB of RAM, running Debian GNU/Linux 12, this error occurred when calculating the 12th-order derivative of the function $f(x)=\sin(x)\,\exp(-x^{2})$, composed with itself five times.
-In contrast, ForwardDiff, implemented in Julia, computes derivatives using dual numbers and elegantly handles higher-order derivatives through their recursive nesting.
-For simplicity, we refer to a structure resulting from nesting these numbers $n$ times as a dual number of order $n$ or a multidual number of order $n$ [16, 17, 18, 19].
-Despite the elegance of this approach, challenges emerge when computing higher-order derivatives with ForwardDiff.
-With each successive derivative, ForwardDiff generates a new Dual type that encapsulates all previously computed derivative information.
-This recursive nesting increases complexity, as each successive derivative introduces additional layers of dual numbers.
-For higher-order derivatives, this nesting becomes particularly problematic, leading to excessive memory consumption and ultimately causing stack overflows.
-In our numerical experiments, a StackOverflowError occurred when attempting to compute the 15th-order derivative of the previously mentioned function.
-In this study, we present a Fortran implementation of dual numbers to arbitrary order without relying on a recursive or nested structure.
-This approach theoretically enables the computation of derivatives of any order, effectively overcoming the limitations of the nested dual number method.
-While practical constraints, such as computational resources, will ultimately limit the achievable order of derivatives, our method significantly extends the range of computable derivatives beyond what is feasible with the nested approach.
-It is important to note, however, that as the derivative order increases, the risk of precision loss grows.
-Despite depending on the specific function being differentiated, our numerical experiments show that derivatives of orders around 100 may require quadruple precision for accurate computation.
-For even higher-order derivatives, further increases in precision will likely be necessary.
-Our Fortran implementation uses complex numbers as the components of dual numbers.
-Although an implementation with real numbers may be necessary for efficiency reasons, we chose to use complex numbers for greater generality.
-It is important to note the lack of a consistent convention regarding the branches chosen for complex functions across different programming environments.
-For example, when computing $\arcsin(1.1)$, the result obtained from SBCL (version 2.2.9.debian), Octave (version 7.3.0), and Wolfram Mathematica (version 12.0) is $(1.5708,-0.44357)$.
-In contrast, the same computation using Python’s cmath library yields $(1.5708,0.44357)$, a result consistent with that produced by modern versions of Fortran.
-We have elected to adopt the latter convention, aligning with the Fortran implementation, to avoid the necessity of developing custom functions to select a specific branch.
-The decision regarding which branch to standardize will ultimately rest with the Fortran standards committee (ISO/IEC JTC1/SC22/WG5).
+Differentiation is a fundamental operation in science and engineering, with applications ranging from optimization and sensitivity analysis to numerical methods and scientific modeling.
+Derivatives can be computed symbolically using computer algebra systems such as Maxima, Maple, or Mathematica, or numerically through finite-difference schemes.
+While symbolic differentiation provides exact expressions and finite differences are straightforward to implement, both approaches have well-known limitations related to expression growth, numerical stability, and truncation errors.
+An alternative approach is Automatic Differentiation (AD) [9, 4, 2], which computes derivatives with machine precision by systematically applying the chain rule to numerical programs, without resorting to symbolic manipulation.
+Automatic Differentiation is an algorithmic technique that enables the efficient computation of derivatives of functions defined by computer programs.
+Although AD has been extensively studied, most practical implementations focus on first- and second-order derivatives of real-valued functions [24, 23, 11, 12, 30, 8, 26].
+One notable implementation is ADOL-C [10, 29, 1], which employs operator overloading to support differentiation in C/C++.
+While ADOL-C provides partial support for higher-order derivatives, several elementary functions, including inverse trigonometric functions and exponentiation, are not fully supported across all derivative orders.
+Several modern AD frameworks provide mechanisms for computing higher-order derivatives of arbitrary order.
+Graph-based approaches, such as those used in the PyTorch library [17], rely on dynamic computational graphs that record intermediate operations at runtime.
+Although this strategy enables repeated differentiation, it requires retaining the entire computational graph in memory, leading to rapid memory growth as the derivative order increases.
+In practice, this often results in memory exhaustion when computing higher-order derivatives of moderately complex functions.
+An alternative strategy is based on forward-mode AD using dual numbers.
+The ForwardDiff package in Julia [25] supports higher-order derivatives by recursively nesting dual numbers.
+For simplicity, we refer to a structure obtained by nesting dual numbers $n$ times as a dual number of order $n$, or a multidual number of order $n$ [22, 21, 15, 28].
+While this approach is mathematically elegant, the recursive nesting leads to rapid growth in memory usage and computational complexity.
+As the derivative order increases, the nested structure becomes increasingly inefficient and may ultimately result in stack overflows or memory exhaustion.
+In this work, we present a Fortran-based implementation of dual numbers that supports the computation of derivatives of arbitrary order without relying on recursive or nested data structures.
+By employing a direct representation of dual numbers, the proposed approach avoids the memory explosion inherent to nested methods and significantly extends the range of derivative orders that can be computed in practice.
+While practical limits are ultimately imposed by available computational resources and numerical precision, the proposed method overcomes key scalability limitations of existing approaches and is particularly well suited for high-performance scientific computing applications.
+To the authors’ knowledge, this work presents the first practical implementation of dual numbers of arbitrary order that avoids recursive or nested representations while remaining scalable for very high derivative orders.
 2 Dual numbers and derivatives
-2.1 First order case
-Analogous to the definition of a complex number $z=a+i\,b$ where $a$, $b$ are real numbers and $i$, the imaginary unit satisfying $i^{2}=-1$, a dual numbers is defined as
-$r$ $=a_{0}\,\epsilon_{0}+\epsilon_{1}\,a_{1},$ (1)
-$r$ $=a_{0}+\epsilon_{1}\,a_{1},$ (2)
-where $a_{0}$, $a_{1}$ are real or complex numbers, $\epsilon_{0}=1$, and $\epsilon_{1}$ is the dual unit satisfying
+2.1 First-order case
+Analogous to the definition of a complex number z=a+ibz=a+i\,b, where $a,b\in\mathbb{R}$ and $i$ is the imaginary unit satisfying $i^{2}=-1$, a dual number is defined as
+$r$ $=a_{0}\,\epsilon_{0}+a_{1}\,\epsilon_{1},$ (1)
+$=a_{0}+a_{1}\,\epsilon_{1},$ (2)
+where $a_{0}$ and $a_{1}$ are real or complex numbers, $\epsilon_{0}=1$, and $\epsilon_{1}$ is the dual unit satisfying
 $\epsilon_{1}^{2}=0.$ (3)
-In a similar manner to how an analytical real function is extended to a complex function, we can evaluate the Taylor expansion of a function $f(z)$ at $z+\epsilon_{1}$:
+In a manner analogous to extending a real function to the complex domain, an analytic function $f(z)$ can be evaluated at a dual argument $z+\epsilon_{1}$ by means of its Taylor expansion,
 $f(z+\epsilon_{1})=f(z)+f^{\prime}(z)\,\epsilon_{1}+\frac{1}{2}f^{\prime\prime}(z)\,\epsilon_{1}^{2}+\cdots.$ (4)
-However, in virtue of Eq. (4) $\epsilon_{1}^{k}=0$ for $k>1$. This simplifies the expansion to
+However, due to Eq. (3), all powers $\epsilon_{1}^{k}$ vanish for $k>1$, and the expansion reduces to
 $f(z+\epsilon_{1})=f(z)+f^{\prime}(z)\,\epsilon_{1}.$ (5)
-Therefore, evaluating an analytic function at the dual number $z+\epsilon_{1}$ yields a dual number where its $\epsilon_{0}$ component is $f(z)$ and its $\epsilon_{1}$ component is $f^{\prime}(z)$.
-The process of extending $f(z)$ to operate on dual numbers is called dualizing $f(z)$.
-As a simple example, the dual extension of the sine function $sin(z)$ will be
-$sin(z)=\sin z+\cos z\;\epsilon_{1}.$ (7)
-Here, the function name in italics indicates the dual version of the original function.
-It is important to note that this function is not the general dual version for the “$\sin$” function, as $z$ is a complex number.
-To construct the general dual version we need to be able of computing $sin(g)$ for $g=g_{0}+\epsilon_{1}\,g_{1}$ a dual number.
-Setting $g(z+\epsilon_{1})$ = $g_{0}+g_{1}\epsilon_{1}$ and evaluating in the Taylor series expansion for the function $f$, we have
-$f(g_{0}+\epsilon_{1}\,g_{1})=f(g_{0})+f^{\prime}(g_{0})g_{1}\,\epsilon_{1}.$ (8)
-Therefore, the dual extension of the sine function becomes:
-$sin(g)=\sin(g_{0})+\cos(g_{0})\,g_{1}\,\epsilon_{1}.$ (9)
-This procedure can be applied to dualize all elementary functions [^1] A non-elementary function, such as the error function $\text{erf}(x)$, can be extended to dual numbers in the same way.
-However, since this function is not implemented in Fortran for complex values, it is not extended in our dual number implementation.
-and main operators of a programming language.
-The generalization for calculating arbitrary-order derivatives is straightforward from a theoretical standpoint.
-However, computational implementation presents significant challenges, particularly in avoiding stack overflow errors that can occur at higher derivative orders when using recursive or nested structures.
-Most works in the literature focus on implementing dual numbers for the real case, typically up to the first or second order [20, 21, 22, 23, 19].
-Some extensions to third and fourth orders, including the complex case, have also been proposed [16, 17].
-While implementations using recursive or nested structures are effective and efficient for lower-order derivatives, they may encounter limitations at very high derivative orders due to stack depth limitations.
-Achieving dual number implementations for arbitrary orders without relying on recursion is the primary contribution of our study.
-2.2 arbitrary order case
-Derivatives to arbitrary order can be calculated by defining the dual number [17]
+Thus, evaluating an analytic function at the dual number $z+\epsilon_{1}$ produces a dual number whose $\epsilon_{0}$ component is $f(z)$ and whose $\epsilon_{1}$ component is $f^{\prime}(z)$.
+The extension of $f(z)$ to operate on dual numbers is commonly referred to as dualizing the function.
+As a simple example, the dual extension of the sine function evaluated at $z+\epsilon_{1}$ is
+$\sin(z+\epsilon_{1})=\sin z+\cos z\,\epsilon_{1}.$ (6)
+Here, the italicized function name denotes the dual version of the original function.
+This expression, however, corresponds only to the special case in which the argument is $z+\epsilon_{1}$, with $z$ a complex number.
+To construct the general dual extension, consider a dual number $g=g_{0}+g_{1}\,\epsilon_{1}$.
+Substituting this expression into the Taylor expansion of an analytic function $f$, we obtain
+$f(g_{0}+g_{1}\,\epsilon_{1})=f(g_{0})+f^{\prime}(g_{0})\,g_{1}\,\epsilon_{1}.$ (7)
+Accordingly, the general dual extension of the sine function is given by
+$\sin(g)=\sin(g_{0})+\cos(g_{0})\,g_{1}\,\epsilon_{1}.$ (8)
+This procedure can be applied to dualize all elementary functions and the main operators of a programming language [^1] Non-elementary functions, such as the error function $\mathrm{erf}(x)$, can in principle be extended to dual numbers in the same manner.
+However, since this function is not implemented for complex arguments in Fortran, it is not included in the present implementation..
+From a theoretical perspective, the generalization of this approach to higher-order derivatives is straightforward.
+In practice, however, computational implementation becomes increasingly challenging, particularly when recursive or nested data structures are employed.
+Most existing implementations focus on real-valued dual numbers and are restricted to first- or second-order derivatives [5, 7, 31, 19, 28].
+Extensions to third- and fourth-order derivatives, including the complex case, have also been proposed [22, 21].
+While recursive and nested representations are effective for low derivative orders, they may encounter severe limitations at higher orders due to excessive memory usage and stack depth constraints.
+Eliminating these limitations by avoiding recursion is a central motivation of the present work.
+2.2 Arbitrary-order case
+Derivatives of arbitrary order can be computed by defining a dual number of order $n$ as [21]
 $r_{n}=\sum_{k=0}^{n}a_{k}\,\epsilon_{k},$ (9)
-with $a_{k}$ a complex number and
-$\epsilon_{i}\cdot\epsilon_{j}=\begin{cases}0&\mbox{if }i+j>n,\vspace*{0.3cm}\\ \frac{(i+j)!}{i!\,j!}\epsilon_{i+j}&\mbox{otherwise. }\end{cases}$ (11)
-Now, evaluating the Taylor expansion for an analytical function in the number $z+\epsilon_{1}$ and using Eq.(11) we have
-$f(z+\epsilon_{1})=f(z)\epsilon_{0}+f^{\prime}(z)\epsilon_{1}+\dots+f^{(n)}(z)\epsilon_{n}.$ (12)
-Computing Eq.
-(12) for the elementary function is not very complicated, and thus the dual extension for $f(z)$ is relatively easy.
-However, in the general case, the dual version for $f(g(z))$, i.e.
-$f(g)$ is much more complicated.
-Nevertheless, the process is simplified if one first implements the well known Faà Di Bruno formula [24, 25] for dual numbers as it is showed in the next section.
+where $a_{k}$ are complex coefficients and the basis elements $\epsilon_{k}$ satisfy the multiplication rule
+$\epsilon_{i}\cdot\epsilon_{j}=\begin{cases}0,&\text{if }i+j>n,\\[5.69046pt] \dfrac{(i+j)!}{i!\,j!}\,\epsilon_{i+j},&\text{otherwise}.\end{cases}$ (10)
+Evaluating the Taylor expansion of an analytic function at $z+\epsilon_{1}$ and using Eq. (10) yields
+$f(z+\epsilon_{1})=f(z)\,\epsilon_{0}+f^{\prime}(z)\,\epsilon_{1}+\cdots+f^{(n)}(z)\,\epsilon_{n}.$ (11)
+The dual extension of elementary functions can therefore be constructed directly.
+In the general case of a composite function $f(g(z))$, the corresponding dual extension $f(g)$ becomes significantly more involved.
+This difficulty can be addressed by implementing the Faà di Bruno formula [13, 6] within the dual number framework, as discussed in the following section.
 3 Dual number implementation to arbitrary order
-To implement dual numbers to arbitrary order, we will first implement the chain rule to calculate the $n$-th derivative of a function in Fortran. The data structure we will use is shown in Listing 1.
-1 type, public :: dualzn
-2 complex(prec), allocatable, dimension(:) :: f
-3 end type dualzn
-Listing 1: Fortran Code for dualzn Type using the precision given by prec. The $k$-th component of a dual number $g$ of this type is accessed by $g\%f(k)$.
-Now, according to the famous Faà Di Bruno formula, the $n$-th derivative of $f(g(x))$ is given by
+To implement dual numbers of arbitrary order, we first construct a non-recursive formulation of the chain rule suitable for numerical evaluation in Fortran.
+The core data structure employed throughout this work is the derived type shown in Listing 1, which stores the coefficients of a dual number of order $n$ in a one-dimensional array.
+1type, public :: dualzn
+2 complex(prec), allocatable :: f(:)
+3end type dualzn
+Listing 1: Fortran derived type dualzn using the precision specified by prec. The $k$-th coefficient of a dual number $g$ is accessed as g%f(k).
+This representation avoids recursive or nested data structures and allows all derivative components to be accessed directly by index, which is essential for scalability at high derivative orders.
+3.1 Chain rule via Faà di Bruno formula
+According to the Faà di Bruno formula, the $n$-th derivative of a composite function $f(g(x))$ is given by
 $D^{n}f(g(x))=\sum\frac{n!}{k_{1}!k_{2}!\cdots k_{n}!}f^{(k_{1}+k_{2}+\cdots+k_{n})}(g(x))\prod_{j=1}^{n}\left(\frac{g^{(j)}(x)}{j!}\right)^{k_{j}},$ (12)
-where $D^{n}=d^{n}/dx^{n}$, $f^{(k)}(x)=D^{k}f(x)$ and the the sum is taken over all sequences of nonnegative integers $k_{1},k_{2},\ldots,k_{n}$ such that they obey the Diophantine equation:
+where $D^{n}=d^{n}/dx^{n}$, $f^{(k)}(x)=D^{k}f(x)$, and the sum is taken over all nonnegative integer solutions of the Diophantine equation
 $k_{1}+2k_{2}+3k_{3}+\cdots+nk_{n}=n.$ (13)
-Although a recursive formula that avoid solving the Diophantine equation can be obtained [26], a more direct approach is to write Eq. (13) as:
-$D^{n}f(g(x))=\sum_{k=1}^{n}f^{(k)}(g(x))\;B_{n,k}(g^{\prime}(x),g^{\prime\prime}(x),\dots,g^{(n-k+1)}(x)),$ (15)
-where the $B_{n,k}(x_{1},x_{2},\dots,x_{n-k+1})$ are the partial Bell polynomials which can be recursively defined by [27]
-$B_{n,k}(x_{1},x_{2},\dots,x_{n-k+1})=\sum_{i=0}^{n-k}\binom{n-1}{i}x_{i+1}B_{n-i-1,k-1}(x_{1},x_{2},\dots,x_{n-k-i+1}),$ (16)
-with
-$B_{0,0}(\cdot)$ $=1$ (17)
-$B_{n,0}(\cdot)$ $=0\text{~{}for~{}}n\geqslant 1$ (18)
-$B_{0,k}(\cdot)$ $=0\text{~{}for~{}}k\geqslant 1.$ (19)
-From this recursive definition, we can build the iterative version shown in Algorithm 3.
+Although recursive formulations that avoid explicitly solving this Diophantine equation exist [16], a more convenient and computationally efficient formulation is obtained by rewriting Eq. (12) as
+$D^{n}f(g(x))=\sum_{k=1}^{n}f^{(k)}(g(x))\,B_{n,k}\!\left(g^{\prime}(x),g^{\prime\prime}(x),\dots,g^{(n-k+1)}(x)\right),$ (14)
+where $B_{n,k}$ are the partial Bell polynomials.
+The partial Bell polynomials can be defined recursively as [3]
+$B_{n,k}(x_{1},\dots,x_{n-k+1})$ $=\sum_{i=0}^{n-k}\binom{n-1}{i}\,x_{i+1}\,B_{n-i-1,k-1}(x_{1},\dots,x_{n-k-i+1}),$ (15)
+with initial conditions
+$B_{0,0}$ $=1,$ (16)
+$B_{n,0}$ $=0\quad(n\geq 1),$ (17)
+$B_{0,k}$ $=0\quad(k\geq 1).$ (18)
+From this definition, an iterative dynamic-programming implementation can be constructed. Algorithm 3.1 presents pseudocode for the evaluation of $B_{n,k}$.
 […]
-1:$n,k\in\mathbb{Z}$, $x\in\mathbb{C}^{m}$
-2:$result\_value\in\mathbb{C}$
-3:Initialize $dp[n+1,k+1]\leftarrow 0$
-4:$dp[1,1]\leftarrow 1$
+1 Iterative computation of the partial Bell polynomial $B_{n,k}(x)$.
+[…]
+[…]
+3:Initialize $dp[0:n,0:k]\leftarrow 0$
+[…]
 5:for $nn\leftarrow 1$ to $n$ do
-6:     $dp[nn+1,1]\leftarrow 0$
-7:end for
-8:for $kk\leftarrow 1$ to $k$ do
-9:     $dp[1,kk+1]\leftarrow 0$
+6:  for $kk\leftarrow 1$ to $\min(nn,k)$ do
+7:   for $i\leftarrow 0$ to nnkknn-$k$ do
+8:     $dp[nn,kk]\leftarrow dp[nn,kk]+\binom{nn-1}{i}\,x_{i+1}\,dp[nn-i-1,kk-1]$
+9:   end for
 10:  end for
-11:if $n=0$ and $k=0$ then
-12:     $result\_value\leftarrow 1$
-13:     return $result\_value$
-14:else if $\text{size}(x)=0$ then
-15:     $result\_value\leftarrow 0$
-16:     return $result\_value$
-17:end if
-18:for $nn\leftarrow 1$ to $n$ do
-19:     for $kk\leftarrow 1$ to $k$ do
-20:         $LX\leftarrow\text{size}(x)$
-21:         $nx\leftarrow\max(nn-kk+1,LX)$
-22:         if $nx>0$ then
-23:              Allocate $newx[nx]\leftarrow 0$
-24:              $newx[1:LX]\leftarrow x$
-25:              $sum\_val\leftarrow 0$
-26:              for $ii\leftarrow 0$ to $nn-kk$ do
-27:                  $sum\_val\leftarrow sum\_val+\binom{nn-1}{ii}\cdot newx[ii+1]\cdot dp[nn-ii,kk]$
-28:              end for
-29:              $dp[nn+1,kk+1]\leftarrow sum\_val$
-30:              Deallocate $newx$
-31:         end if
-32:     end for
-33:end for
-34:$result\_value\leftarrow dp[n+1,k+1]$
-35:return $result\_value$
-Having implemented the $B_{n,k}$ function, we can now implement the chain rule Eq.
-(15) in Fortran using the derived data type dualzn from Listing 1 and following Algorithm 3.
-Here, the $fci$ function provides the dual version of $f(z)$.
-We emphasize that it is not the dual version of $f(g(z))$, which is constructed after applying the chain rule to $f$ through function Dnd.
-Let us exemplify this with the sinusoidal function.
+11:end for
+12:return $dp[n,k]$
+3.2 Implementation of the chain rule
+Once the partial Bell polynomials are available, the chain rule (14) can be implemented directly for dual numbers of type dualzn.
+Algorithm 3.2 shows the pseudocode for the function Dnd, which computes the $n$-th derivative of $f(g(x))$.
+The function fci provides the dual extension of $f(z)$ evaluated at a scalar argument.
 […]
-1:$fci$: Function, $gdual$: dualzn type, $n\in\mathbb{Z}$
-2:$dnfcg\in\mathbb{C}$
-3:$g0\leftarrow gdual\%f(0)$
-4:$fvd\leftarrow fci(g0)$
+2 Pseudocode for Dnd(fci, gdual, n) implementing the chain rule.
+1:fci: function, gdual: dualzn, $n\in\mathbb{Z}$
+[…]
+[…]
+[…]
 5:if $n=0$ then
-6:     $dnfcg\leftarrow fvd\%f(0)$
-7:else
-8:     $suma\leftarrow 0$
+6:  return $fvd\%f(0)$
+7:end if
+[…]
 9:for $k\leftarrow 1$ to $n$ do
-10:         Allocate $xvg[1:n-k+1]$
-11:         for $j\leftarrow 1$ to $n-k+1$ do
-12:              $xvg[j]\leftarrow gdual\%f(j)$
-13:         end for
-14:         $suma\leftarrow suma+fvd\%f(k)\cdot B(n,k,xvg)$
-15:         Deallocate $xvg$
-16:     end for
-17:     $dnfcg\leftarrow suma$
-18:end if
-19:return $dnfcg$
-The dual version $sin(z)$ to order $n$, of the $\sin(z)$ function is given by
-$sin(z)$ $=\sum_{k=0}^{n}D^{k}(\sin z)\,\epsilon_{k}$
-$sin(z)$ $=\sum_{k=0}^{n}\sin\left(z+k\,\frac{\pi}{2}\right)\epsilon_{k}.$ (20)
-We recall that in this function $sin\;z$ is a dual number but $z$ is not. To construct the dual version $sin(g)$, being $g$ a dual number, we need to apply the Dnd function to $sin\;z$. Therefore
-$sin(g)=\sum_{k=0}^{n}\text{Dnd}(sin,g,k)\,\epsilon_{k}$ (21)
-will be the dual version of $\sin(g(z))$.
-By following this procedure, the dual version for all elementary functions can be obtained.
-As exemplified in Eq.
-(20), to obtain the dual version for $f(z)$ its derivatives are necessary.
-For most of the elementary functions this can be done with no much effort.
-There are another functions, however, for which to obtain its dual version $f(z)$ could be a little more complicated.
-Take for instance the inverse of the sine function, the $\arcsin(z)$ function.
-A recursive [^2] While a recursive approach to calculating the derivatives of functions may appear more direct, we advise against it due to potential inefficiencies, including lower performance and higher memory consumption.
-formula to calculate its $n$-th derivative is given by
-$D^{0}(\arcsin z)$ $=\arcsin z$ (22)
-$D^{1}(\arcsin z)$ $=\frac{1}{\sqrt{1-z^{2}}}$ (23)
-$D^{n}(\arcsin z)$ $=-\sum_{k=1}^{n-1}\binom{n-1}{k}D^{k}\left(\sqrt{1-z^{2}}\right)\frac{D^{n-k}(\arcsin z)}{\sqrt{1-z^{2}}}.$ (24)
-One can pursue this approach; however, if we first dualize the functions $inv(z)=1/z$ as well as the square root function and the difference and multiplication operators, then by expressing Eq.
-(23) in its dual form, all higher-order derivatives can be automatically computed.
-All these function are relatively easy to dualize, for instance, Algorithms 3 and 4 shows the dualization of the multiplication operator using the Leibniz rule.
-Algorithm 3 Pseudocode for the function timesdzn(A,B,k), the Leibniz rule.
-1:$A$: dualzn type, $B$: dualzn type, $k\in\mathbb{Z}$
-2:$fr\in\mathbb{C}$
-3:$fr\leftarrow 0$
+10:  $sum\leftarrow sum+fvd\%f(k)\,B_{n,k}(gdual\%f(1:n-k+1))$
+11:end for
+12:return sumsum
+3.3 Dualization of elementary functions
+As an illustrative example, the dual extension of $\sin(z)$ to order $n$ is
+$\sin(z)$ $=\sum_{k=0}^{n}D^{k}(\sin z)\,\epsilon_{k}=\sum_{k=0}^{n}\sin\!\left(z+k\frac{\pi}{2}\right)\epsilon_{k}.$ (19)
+Here, $\sin(z)$ denotes a dual number whose argument $z$ is scalar. To construct the dual extension $\sin(g)$, where $g$ is a dual number, the chain rule must be applied:
+$\sin(g)=\sum_{k=0}^{n}\texttt{Dnd}(\sin,g,k)\,\epsilon_{k}.$ (20)
+Following this procedure, dual extensions for all elementary functions can be constructed.
+In cases where closed-form expressions for higher derivatives are cumbersome, functions can be dualized by combining previously dualized elementary operations.
+For example, although the higher derivatives of $\arcsin(z)$ satisfy the recursive relations
+$D^{0}(\arcsin z)$ $=\arcsin z,$ (21)
+$D^{1}(\arcsin z)$ $=\frac{1}{\sqrt{1-z^{2}}},$ (22)
+$D^{n}(\arcsin z)$ $=-\sum_{k=1}^{n-1}\binom{n-1}{k}D^{k}\!\left(\sqrt{1-z^{2}}\right)\frac{D^{n-k}(\arcsin z)}{\sqrt{1-z^{2}}},$ (23)
+it is computationally preferable to dualize the inverse, square root, multiplication, and subtraction operators, and to rely on the automatic propagation of derivatives through these operations.
+As an example, Algorithms 3 and 4 show the dualization of multiplication using the Leibniz rule.
+Algorithm 3 Leibniz rule for the $k$-th derivative of a product.
+1:$A,B$: dualzn, $k\in\mathbb{Z}$
+[…]
+[…]
 4:for $i\leftarrow 0$ to $k$ do
-5:     $fr\leftarrow fr+\binom{k}{i}\cdot A\%f(i)\cdot B\%f(k-i)$
+5:  $res\leftarrow res+\binom{k}{i}A\%f(i)B\%f(k-i)$
 6:end for
-7:return $fr$
-Algorithm 4 Pseudocode for the function timesd(A,B) which implements the product of the dual numbers A and B
-1:$A$: dualzn type, $B$: dualzn type
-2:$fr$: dualzn type
-3:Allocate $fr\%f(0:\text{order})$
+7:return resres
+Algorithm 4 Product of two dual numbers.
+1:$A,B$: dualzn
+2:ABAB: dualzn
+3:Allocate $res\%f(0:\texttt{order})$
 4:for $k\leftarrow 0$ to order do
-5:     $fr\%f(k)\leftarrow\texttt{timesdzn}(A,B,k)$
+5:  $res\%f(k)\leftarrow\texttt{timesdzn}(A,B,k)$
 6:end for
-7:return $fr$
+7:return resres
 4 The DNAOAD package
-With the just presented theory, dual numbers to arbitrary order are implemented in the so-called DNAOAD package.
-This implementation not only allows for computation of derivatives also to arbitrary order, but also allows the construction of algorithms in the context of dual numbers.
-Before presenting some interesting applications we first describe the DNAOAD package.
+Building on the theory introduced in the previous sections, we present DNAOAD, a Fortran implementation of dual numbers of arbitrary order.
+In addition to enabling the computation of derivatives of arbitrary order, DNAOAD allows users to formulate numerical algorithms directly in the algebra of dual numbers, so that derivatives are propagated automatically through standard arithmetic operations and elementary functions.
+This section describes the structure and distribution of the package before presenting representative applications.
 4.1 Elements of the package
-The implementation is available for double (real64) and quadruple (real128) precisions.
-The user can choose which precision to use.
-The module for double precision is shown in Listing 2.
-This module is used by the dualzn_mod module, which is the core of the dual number implementation and contains approximately 2,000 lines of code.
-The first 28 lines, which also show the functions and operators extended to dual numbers (overloaded), are shown in Listing 3.
-1module precision_mod
-2 use iso_fortran_env, only: int32, int64, real32, real64, real128
-3 implicit none
-4
-5 integer, parameter :: prec = real64
-6end module precision_mod
-Listing 2: The precision module for double precision implementation.
+DNAOAD is distributed in two equivalent forms.
+The first is a standalone source distribution intended for traditional workflows, such as building with make or integrating the source files directly into an existing Fortran code base [20].
+The second is an FPM (Fortran Package Manager) distribution [18], which automatically handles compilation and linking and therefore provides a more convenient and reproducible workflow.
+While both distributions offer identical functionality, the FPM-based interface is recommended for new projects and for rapid integration into larger Fortran applications.
+The core of the DNAOAD package is provided by the module dualzn_mod, which defines the dualzn derived type and implements the overloaded operators and intrinsic-like functions required for seamless manipulation of dual numbers.
+For illustration, the first 28 lines of dualzn_mod are shown in Listing 2.
 1module dualzn_mod
@@
 28 public :: operator(**), operator(==), operator(/=)
-Listing 3: Extract from the module dualzn_mod
-The description of the variables, functions and operators in listing 3 is as follows.
+Listing 2: Extract from the module dualzn_mod
+The description of the variables, functions and operators in listing 2 is as follows.
   1. 1.
@@
 4.2 Usage of DNAOAD
-The package is available for Windows and GNU/Linux in both double and quadruple precision, and it supports the gfortran and ifx Intel Fortran compilers.
-It can be downloaded from [30].
-After uncompressing the file, the user can navigate, for instance, to the folder:
-*/DNAOAD/GNU_Linux/ifx_case/prec_real64/examples/’ (the ‘*’ depends on where the package is uncompressed on the local machine), and run some of the examples.
-The source code is also provided; however, in our examples, we will assume that the precompiled libraries in the GNU_Linux folder are used.
+The DNAOAD package is available for both Windows and GNU/Linux platforms in double and quadruple precision, and it supports the gfortran and Intel ifx Fortran compilers.
+The source code and precompiled libraries can be obtained from the project repository [20].
+The distribution includes a set of example programs that illustrate the basic usage of the library.
+For the standalone distribution, users may compile and run the example codes using either the provided precompiled libraries or by building the sources directly.
+In the examples presented here, we assume that the precompiled libraries supplied with the package are used.
+Detailed build instructions and platform-specific scripts are included in the distribution to simplify compilation on both GNU/Linux and Windows systems.
 4.2.1 Simple examples
-As a simple example, the program in listing 4 computes the derivatives from the 0-th up to the 5-th order of $f=\sin(z)^{\log(z^{2})}$ at $z_{0}=1.1+2.2i$.
-Assuming the code is saved in the folder ‘*/DNAOAD/GNU_Linux/ifx_case/prec_real64/examples/’ under the file example1.f90, it can be compiled in GNU/Linux as:
-ifx -I../LibDualzn64 -o example1.out example1.f90 -L../LibDualzn64 -ldualzn
-and in Windows as:
-ifx -o example1.exe example1.f90 -I../LibDualzn64 ../LibDualzn64/libdualzn.lib
-For the gfortran compiler, use:
-gfortran -I../LibDualzn64 -o example1.exe example1.f90 -L../LibDualzn64 -ldualzn
-Alternatively, run the provided script create_exec.sh (on GNU/Linux) or create_exec.bat (on Windows), followed by the file name where the code is saved.
+As a simple example, Listing 3 shows a program that computes the derivatives from zeroth to fifth order of the function $f(z)=\sin(z)^{\log(z^{2})}$ evaluated at $z_{0}=1.1+2.2\,i$.
+When using the standalone distribution, the program can be compiled by linking against the DNAOAD library with either the ifx or gfortran compilers, following the instructions provided with the package.
+For convenience, platform-specific scripts are also included to automate the compilation process.
+When using the FPM (Fortran Package Manager) distribution, compilation and execution are fully managed by FPM.
+In this case, the example can be built and executed with a single command, for instance, fpm run ex1, which significantly simplifies the workflow.
 1program main
@@
 26end program main
-Listing 4: Example of derivative calculation. Since a dualzn number is an allocatable entity, it must first be initialized.
+Listing 3: Example of derivative calculation. Since a dualzn number is an allocatable entity, it must first be initialized.
 Dual numbers can also be used to differentiate Fortran code.
@@
 This calculation is virtually impossible symbolically, and finite differences would be inefficient and inaccurate.
-Listing 5 shows a program to compute these derivatives.
-Appendix A., presents Python and Julia versions when the funcion is nested 5 times and the order of derivation is 10.
+Listing 4 shows a program to compute these derivatives.
+Appendix A presents Python and Julia versions when the funcion is nested 5 times and the order of derivation is 10.
 While this example is theoretical, real-world scenarios often involve multiple function compositions, vector rotations, and similar operations.
@@
 47end program main
-Listing 5: Example of differentiating a function that is not given in closed form but implemented as a computer program.
+Listing 4: Example of differentiating a function that is not given in closed form but implemented as a computer program.
 4.2.2 Calculating gradients, Jacobians and Hessians
@@
 Below an example of use of this module.
-This example can be found in the folder:
-*/DNAOAD/GNU_Linux/ifx_case/prec_real64/examples/’.
 1!module with example of functions
@@
 47 integer, parameter :: nf =3, mq = 4
-48 complex(prec), parameter :: ii = (0,1)
+48 complex(prec), parameter :: $i$ = (0,1)
 49 complex(prec), dimension(mq) :: q, vec
@@
 56 vec = [1.0_prec,2.0_prec,3.0_prec,4.0_prec]
-57 q = vec/10.0_prec + ii
+57 q = vec/10.0_prec + $i$
 58
@@
 87end program main
-Listing 6: Example of using the diff_mod.
+Listing 5: Example of using the diff_mod.
 5 Conclusions
-This study addresses the challenges of computing higher-order derivatives with dual numbers by introducing DNAOAD, a Fortran-based implementation designed for arbitrary-order automatic differentiation.
-Most existing dual-number implementations are limited to lower-order derivatives, with only one notable approach—Julia’s ForwardDiff—supporting arbitrary-order differentiation through nested dual numbers.
-However, no prior work has demonstrated a method for arbitrary-order differentiation with dual numbers that avoids reliance on nested structures, which are prone to significant memory consumption and stack overflow errors.
-DNAOAD bridges this gap by directly implementing dual numbers to arbitrary orders, eliminating the need for recursive or nested frameworks.
-This direct and innovative approach significantly reduces computational overhead, enabling the efficient and reliable computation of derivatives at very high orders, and providing a scalable and practical solution to the limitations of existing methods.
-Through numerical experiments, DNAOAD demonstrated its capability to compute derivatives at very high orders with remarkable accuracy, though extremely high orders may require enhanced numerical precision, such as quadruple precision.
-The implementation also provides a considerable number of useful functions for various applications, including the calculation of gradients, Jacobians, and Hessians, as well as higher-order derivatives of complex functions.
-These capabilities enhance DNAOAD’s applicability to a broad range of fields, such as optimization, physics, and engineering, where precise differentiation is essential.
-In comparison with existing tools, such as PyTorch and ForwardDiff—which excel in low-order differentiation but face scalability issues due to memory overheads—DNAOAD provides a scalable and efficient alternative.
-Its robustness was demonstrated in experiments involving deeply nested functions, highlighting its potential for tackling complex computational tasks.
-This implementation represents a significant advancement in practical automatic differentiation by combining theoretical rigor with computational efficiency.
-Future work may focus on extending DNAOAD to other programming environments and further optimizing its performance for large-scale applications.
-Data availability statement
-The Fortran source code, which contains the data supporting the findings of this study, is openly available on GitHub at https://github.com/fpenunuri/DNAOAD.
-Appendix Appendix A. Computing higher order derivatives with PyTorch and ForwardDiff
+In this work, we introduced DNAOAD, a Fortran-based implementation of dual numbers designed to support automatic differentiation of arbitrary order.
+Unlike most existing dual-number approaches, which are limited to low-order derivatives or rely on recursive and nested structures, DNAOAD employs a direct, non-nested representation of dual numbers.
+This design avoids the severe memory growth and stack limitations commonly encountered in nested implementations and enables scalable computation of higher-order derivatives.
+The numerical experiments presented in this work indicate that DNAOAD can reliably compute derivatives of substantially higher order than those typically accessible with nested dual-number implementations, without encountering stack overflows or prohibitive memory usage.
+While increasing the derivative order may require enhanced numerical precision—such as quadruple precision for extreme cases—the proposed approach significantly extends the range of derivative orders that can be computed in practice.
+In addition to higher-order derivatives, the implementation provides practical tools for computing gradients, Jacobians, Hessians, and higher-order derivatives of complex-valued functions, making DNAOAD suitable for a wide range of applications in scientific computing, physics, and engineering.
+Compared with existing tools such as PyTorch and ForwardDiff, which perform well for low-order differentiation but face scalability issues due to memory overhead, DNAOAD offers a robust and efficient alternative for high-order differentiation.
+The results highlight the effectiveness of the proposed approach for deeply nested functions and computationally demanding problems.
+Appendix A Computing higher order derivatives with PyTorch and ForwardDiff
 Although the computation of higher-order derivatives can lead to significant memory consumption, the utility of the PyTorch and ForwardDiff libraries remains undeniable.
-Below, we present the example of Listing 5 that demonstrates a relatively small number of function compositions in both Python and Julia, respectively.
+Below, we present the example of Listing 4 that demonstrates a relatively small number of function compositions in both Python and Julia, respectively.
 1import torch
@@
 27print("Elapsed time (s):", end - start)
-Listing 7: Example of computing higher-order derivatives with PyTorch.
+Listing 6: Example of computing higher-order derivatives with PyTorch.
 1using ForwardDiff
@@
 36println("Elapsed time: \$elapsed_time seconds")
-Listing 8: Example of computing higher-order derivatives with ForwardDiff.
+Listing 7: Example of computing higher-order derivatives with ForwardDiff.